# Show Me the World in My Language: Establishing the First Baseline for Scene-Text to Scene-Text Translation

### v2

@@ Abstract
In this work, we study the task of “visually” translating scene text from a source language (e.g., Hindi) to a target language (e.g., English). Visual translation involves not just the recognition and translation of scene text but also the generation of the translated image that preserves visual features of the source scene text, such as font, size, and background. There are several challenges associated with this task, such as translation with limited context, deciding between translation and transliteration, accommodating varying text lengths within fixed spatial boundaries, and preserving the font and background styles of the source scene text in the target language. To address this problem, we make the following contributions: (i) We study visual translation as a standalone problem for the first time in the literature. (ii) We present a cascaded framework for visual translation that combines state-of-the-art modules for scene text recognition, machine translation, and scene text synthesis as a baseline for the task. (iii) We propose a set of task-specific design enhancements to design a variant of the baseline to obtain performance improvements. (iv) Currently, the existing related literature lacks any comprehensive performance evaluation for this novel task. To fill this gap, we introduce several automatic and user-assisted evaluation metrics designed explicitly for evaluating visual translation. Further, we evaluate presented baselines for translating scene text between Hindi and English. Our experiments demonstrate that although we can effectively perform visual translation over a large collection of scene text images, the presented baseline only partially addresses challenges posed by visual translation tasks. We firmly believe that this new task and the limitations of existing models, as reported in this paper, should encourage further research in visual translation.
@@ 1 Introduction
Machine Translation has shown remarkable growth in the last few years, partly attributed to the adoption of neural models [1, 26, 5, 30, 6]. In parallel, substantial advancements have also been made in speech-to-speech translation [20, 10, 15, 12] where the goal is to develop systems that are capable of accurately interpreting spoken language in one dialect and seamlessly translating it into another while preserving the voice of the original speaker, thus enabling effective cross-lingual communication in real-time. Drawing inspiration from these research directions, we present an analogous problem in the scene text domain, namely “scene-text to scene-text translation” or, in short, “visual translation”. The visual translation task aims to translate text present in images from a source language to the target language while preserving the visual characteristics of the text and background as illustrated in Figure 1. Visual Translation has extensive applications, e.g., transforming the travel experience by allowing tourists to instantly understand sign boards in foreign languages and enabling seamless interaction with the visual world without language barriers.
@@
We extensively evaluate proposed baselines for Hindi-to-English and English-to-Hindi visual translation using our newly proposed automatic and user evaluation metrics. Although the proposed baselines shows promise for reasonable visual translation, the overall problem remains far from solved and calls for future research.
We make the following contributions: (i) We study the under-explored task of visual translation that aims to translate text in images to a target language, while preserving its font, style, position, and background. To the best of our knowledge, the comprehensive study of this problem has largely been unexplored in the existing literature. (ii) We introduce a generic cascaded approach for visual translation, and we design a set of baselines using state-of-the-art approaches for scene text recognition, machine translation and scene text synthesis and their task-specific design enhancements. (iii) Training a visual translation model with real-world images is challenging due to the lack of large-scale paired scene text images in different languages. Therefore, we use synthetic images for training. We present a method to generate paired images with words sharing the same visual properties, creating VT-Syn, a parallel corpus of 600K visually diverse English-Hindi word-level translations. To evaluate performance on real images, we provide extensive annotations of translated text from three users. These benchmark datasets will support future research in visual translation. (iv) Due to the lack of principled evaluation metrics for visual translation tasks in the literature, we propose a set of automatic and user evaluation metrics. We believe these metrics will help track the progress of visual translation tasks effectively.
@@ 2 Related Work
Editing Text in Images: The problem of editing text in images has witnessed significant research interest in recent years [27, 28, 24, 16, 14, 23]. This task aims to modify scene text to target content while retaining visual properties of the original text. SRNet [27] is one such method that learns an end-to-end GAN-based style-retention network. SwapText [28] improved upon the SRNet architecture by modeling the geometrical transformation of the text using spatial points. More recently, TextStyleBrush [14], RewriteNet [16], and MOSTEL [23] introduce a self-supervised training approach on real-world images for this task. Further, TextStyleBrush is evaluated on handwritten images as well. Authors in [24] proposed a character-wise text editor model for this task. However, their approach assumes source and target text instances are of the same length, which is not always true, especially in the translation task. A more recent approach MOSTEL [23] also introduces stroke-level modifications to produce more photo-realistic generations. Despite demonstrating a few selected qualitative results for cross-lingual editing, these works have only scratched the surface of this area and it remains largely unexplored and understudied. Our work aims to fill this gap.
@@ 3 Proposed Visual Translation Baseline
Figure 2: Outline of proposed cascaded baseline for Visual Translation. We use state-of-the-art approaches for scene text recognition, machine translation and scene text synthesis to design variants of our baseline. Moreover, we further investigate on scene text synthesis and propose an extension to existing SRNet architecture.
The task of Visual Translation can be reduced to a sequence of sub-tasks: locating and reading text in scene images, translating the text into the target language, and generating the final image containing the translated scene text. Motivated by this observation, we propose a cascaded approach to visual translation by combining models for (i) scene Text Recognition, (ii) machine translation, (iii) scene-text synthesis and (iv) seamlessly blending the generated scene text in the image. These sub-tasks are well-explored independently in computer vision literature; thus, we benefit from the availability of trained models. Further, such an approach can perform generation at the word or phrase level, which can help preserve the consistency of non-text regions in the image. The outline of our cascaded baseline is illustrated in Figure 2. We describe each module in detail in the following sections.
Figure 3: Our baseline variant where we extend scene text synthesis approach SRNet by decoupling the background and foreground generation steps. More details are provided in Section 3.
Training and evaluation of a visual translation baseline requires real-world images in the form of $(I,I^{\prime})$ where $I$ and $I^{\prime}$ are visually identical images containing corresponding scene text in two different languages with matching font and style. However, such instances are not easily available in the real world. We mitigate this challenge of data scarcity by directly synthesizing the desired data: we generate paired images of scene-text which are (i) identical in the image background and are (ii) matching in font and style. A few examples are shown in Figure 4. In generating synthetic samples, we use a large corpus of word translations, as well as a diverse collection of fonts. To simulate real-world scene text, we also render the images on natural backgrounds, as well as vary the orientation, positioning, and size of the scene text in images. A more detailed procedure for generating the synthetic data is provided in Section 4.1.
@@
(ii) Machine Translation of Text. After obtaining the recognized text in the source language $L$, we map each instance to the desired target language $L^{\prime}$ using an off-the-shelf neural machine translation method. We test our model with two state-of-the-art neural machine translation methods, namely IndicTrans2 [7], and M2M100 [6]. IndicTrans2 is trained on a large collection of Indic languages (including Hindi), whereas M2M100 is a more general translation method trained on a diverse collection of languages with support for Indic languages as well.
@@ 3.1 Design Enhancements
To enhance the design, a series of refined and newly introduced steps have been implemented. The process begins with the detection and recognition of text, after which numbers, websites, and email addresses are filtered out using regular expressions. Note that these elements do not need to be translated. Words are then grouped into paragraphs and lines based on the geometry and coordinates of the bounding box, in conjunction with a heuristic function. These paragraphs are translated and segmented into lines, ensuring alignment with the proportion of lines present in each original paragraph. Through cubic spline interpolation, new coordinates for each word within a line are determined, which are then linked back to the original crops of the words. Depending on the new width of the translated words, adjustments are made to the crops—either cutting or replicating them—to maintain the original style of the text.
The process is finalized by accurately positioning the new words on the image with the developed method. Although heuristically designed, this step shows a significant boost in translation quality, as shown in the experiments.
@@ 4.1 VT-Syn: Synthetic Training Data
For training the scene text synthesis components of our pipeline, we need paired images of text in different languages with identical visual properties (style, font, orientation, and size.). It is extremely difficult to get visually identical scene images with text in different languages, and it is even more difficult to generate accurate skeleton images required for training SRNet, MOSTEL, as well as our proposed SRNet++ method. Thus we rely on generating highly diverse synthetic images We introduce VT-Syn, a synthetically generated corpus of 600K visually diverse paired bilingual word images in pairs of English-Hindi as well as Hindi-English.
We utilized an Indic-language scene-text image generator [19] and modified it to generate samples of scene-text in paired languages with controllable parameters for font, style, color, and spatial transformations to ensure visual diversity. Each sample contains source image ($x_{src}$), target word image ($x_{w}$), background image ($x_{bg}$), foreground image ($x_{fg}$) and target image ($x_{tgt}$). We also generate source word images, source and target masks and skeleton images based on the requirements of various scene-text synthesis architectures. We collect 291 publicly available fonts which support both Roman and Devanagari scripts and use a vocabulary of 3K commonly used words in both languages.
A few samples from VT-Syn are shown in Figure 4. Note that the paired image words do not have to be translations as the STS module just has to learn to render the target word with same style as the source image.
Figure 4: VT-Syn dataset examples, which contains paired Eng $\rightarrow$ Hin and Hin $\rightarrow$ Eng images with diverse fonts, text colors, sizes, orientations, and backgrounds (of natural scenes, textures and plain colors).
@@ 4.2 VT-Real: Real Test Dataset
For the purpose of evaluation, we propose VT-Realwhich contains images from ICDAR 2013 [13] and Bharat Scene Text Dataset [11] to evaluate English-to-Hindi and Hindi-to-English translations, respectively. We filter images of moderate complexity [^1] As a first work on Hindi-to-English and English-to-Hindi translation, we have opted not to include highly complex curved and occluded text. from these two sources. In all, our dataset contains 269 images and 1021 words. These images were given to three human annotators to translate the text from Hindi to English and vice versa. A few example translation annotations of this dataset are shown in Figure 5. Even though the above-mentioned datasets have no ground truths for Visual Translation (i.e., scene text in the target language), they are still useful for automatic evaluation proposed in the next section.
Figure 5: A few examples from VT-Real dataset, showing image and Eng-Hin and Hin-Eng ground truth translations, manually annotated by three independent annotators.
@@ 5 Performance Metrics
Evaluating visual translation methods is complex, even more so than evaluating machine translation. While the evaluation of machine translation has been a longstanding research area in the NLP community, the recent research has saturated to use metrics such as BLEU, METOER, ROUGE; visual translation poses several additional challenges. Unlike machine translation which is typically evaluated for a sentence or paragraph of text, in visual translation one has to evaluate correctness of translation for single word or a small set of words or phrase. Further, It requires not only assessing the linguistic accuracy of the translation but also ensuring the preservation of background and font properties.
@@ 5.1 Automatic Evaluation
(iii) $\mathbf{vt\text{-}score}$: For a good quality visual translation, it is important to have a high BLEU score, and perception quality along with font style preservation is required. Due to the absence of a robust automatic model that can verify cross-lingual font style similarity, we only consider Translation Quality and perception quality to compute combined $vt\text{-}score$ as follows:
$vt\text{-}score=\frac{2\cdot TQ\cdot PQ}{TQ+PQ}$ (1)
Please note that, for images that contain only one word, we employ BLEU-1 instead of BLEU-2 to assess translation quality. For the remaining images, BLEU-2 is utilized in the aforementioned scoring computation.
@@ 5.2 User Evaluation
Despite having some automatic evaluation metrics as discussed above, user evaluation is essential for ensuring the accuracy, usability, and effectiveness of visual translations. By involving users, we can gather feedback on clarity of generated translated words, cultural appropriateness, and accessibility. To this end, we propose to have an extensive user evaluation. For performing user evaluation, we hired four human users of age group 20 to 25 who have graduate degree and proficient in both Hindi and English. They were shown the results of each visual translation baseline via Beamer slides. Each slide for metrics (ii) and (iii) featured one output image, while slides for metrics (i) and (iv) displayed both the input and output images together. The proposed user evaluation metrics are described here:
@@
(i) Translation Quality (TQ) (score range: 1-4): This criterion focuses on the accuracy of the translation. Users were asked rate whether the translated text accurately conveys the meaning of the original text. A higher score indicates a more accurate translation. The different ratings by users conveys the following: 4: Linguistically and culturally totally correct translation. 3: Some words are correct, translation can be improved. 2: Very few words are correct and significant improvement required. 1: Totally incorrect translation.
@@
(ii) Readability (R) (score range: 1-4): This criterion evaluates how easily the translated text can be read within the scene image. Factors such as font size, contrast, and placement of the text may influence readability. A higher score indicates better readability. The different readability ratings by users conveys the following: 4: Clearly readable. 3: Can read with some effort. 2: Can read with significant effort, some words not readable. 1: No text present in target language. (iii) Perceptual Quality (PQ) (score range: 1-4): This criterion assesses how well the translated text blends into the scene image, making it difficult to distinguish from a real image. A higher score indicates better integration of the translated text with the scene. Users were asked to rate approaches based on the following: 4: Very clear, looks like real image. 3: Clear image but some patches are there if carefully seen. 2: Lot of patchy effect, looks like a fake image. 1: Too much patchy effect, for sure it is a fake image.
@@
(iv) Source style preservation (SSP) (score range: 1-4): This criterion examines whether the translated text preserves the style, font, color, and other visual attributes of the original text in the scene image. A higher score indicates that the translated text maintains consistency with the source text in terms of visual presentation. 4: Font style, size, color, and background are coherent to the source. 3: Only 2 or 3 of the following: Font style, size, color, and background are coherent to the source. 2: Only 1 or 2 of the following: Font style, size, color, and background are coherent to the source. 1: No source style preservation.
6 Experiments
In this section, we comprehensively evaluate scene text to scene text translation baseline approaches discussed in Section 3.2 using both automatic and user evaluation metrics proposed in Section 5. We use VT-Real introduced in Section 4 for all our evaluation.
The automatic evaluation results are reported in Table 1. We observe that SRNet++ clearly emerges as the best scene text synthesis approach as compared to other existing architectures. The proposed design enhancements also significantly boost translation quality while maintaining nearly identical perceptual quality. We also observe that usage of IndicTrans2 as a translator consistently leads to an increase in translation quality.
The state-of-the-art scene text recognition approaches are as good as ground truth annotations (Oracle) in case of detecting and recognizing English text.
We further perform a rigorous user study using metrics presented in Section 5.2. As discussed in this section, we have collected user feedback from four qualified users and report mean score of TQ, R, PQ and SSP in Table 2. These scores nearly align with observations made via automatic evaluation. We also observe that there is significant room for improvement on all these metrics indicting toward the challenge associated with the task.
Method STR MT STS D.E. TQ (BL-1) TQ (BL-2) PQ VT score
@@ 6 Experiments
B-4 DBNet+ParSeq M2M SRNet++ 3.22 2.19 55.60 3.81
B-3 Oracle M2M SRNet++ 4.20 2.89 55.58 4.97
B-2 Oracle M2M Mostel 2.03 1.40 53.41 2.46
B-1 Oracle M2M SRNet 4.20 2.86 53.82 4.92
Table 1: Automatic Evaluation to evaluate visual translation baseline. We report translation quality (TQ) using BLEU-1 (BL-1) and BLEU-2 (BL-2) metrics and perception quality (PQ). D.E: Design Enhancements. Please refer Section 5.1 for more details.
STR MT STS D.E. TQ R PQ SSP

### v3

@@ Abstract
In this work, we study the task of “visually” translating scene text from a source language (e.g., Hindi) to a target language (e.g., English). Visual translation involves not just the recognition and translation of scene text but also the generation of the translated image that preserves visual features of the source scene text, such as font, size, and background. There are several challenges associated with this task, such as translation with limited context, deciding between translation and transliteration, accommodating varying text lengths within fixed spatial boundaries, and preserving the font and background styles of the source scene text in the target language. To address this problem, we make the following contributions: (i) We study visual translation as a standalone problem for the first time in the literature. (ii) We present a cascaded framework for visual translation that combines state-of-the-art modules for scene text recognition, machine translation, and scene text synthesis as a baseline for the task. (iii) We propose a set of task-specific design enhancements to design a variant of the baseline to obtain performance improvements. (iv) Currently, the existing related literature lacks any comprehensive performance evaluation for this novel task. To fill this gap, we introduce several automatic and user-assisted evaluation metrics designed explicitly for evaluating visual translation. Further, we evaluate presented baselines for translating scene text between Hindi and English. Our experiments demonstrate that although we can effectively perform visual translation over a large collection of scene text images, the presented baseline only partially addresses challenges posed by visual translation tasks. We firmly believe that this new task and the limitations of existing models, as reported in this paper, should encourage further research in visual translation. We have publicly released the code and dataset on our project website: https://vl2g.github.io/projects/visTrans/.
@@ 1 Introduction
Machine Translation has shown remarkable growth in the last few years, partly attributed to the adoption of neural models [2, 30, 6, 34, 7]. In parallel, substantial advancements have also been made in speech-to-speech translation [24, 11, 17, 12] where the goal is to develop systems that are capable of accurately interpreting spoken language in one dialect and seamlessly translating it into another while preserving the voice of the original speaker, thus enabling effective cross-lingual communication in real-time. Drawing inspiration from these research directions, we present an analogous problem in the scene text domain, namely “scene-text to scene-text translation” or, in short, “visual translation”. The visual translation task aims to translate text present in images from a source language to the target language while preserving the visual characteristics of the text and background as illustrated in Fig. 1. Visual Translation has extensive applications, e.g., transforming the travel experience by allowing tourists to instantly understand sign boards in foreign languages and enabling seamless interaction with the visual world without language barriers.
@@
We extensively evaluate the proposed baselines for Hindi-to-English and English-to-Hindi visual translation using our new automatic and user evaluation metrics. While the baselines show promising results, the problem remains far from solved and requires further research.
We make the following contributions: (i) We study the under-explored task of visual translation that aims to translate text in images to a target language while preserving its font, style, position, and background. To the best of our knowledge, the comprehensive study of this problem has largely been unexplored in the existing literature. (ii) We introduce a generic cascaded approach for visual translation, and we design a set of baselines using state-of-the-art approaches for scene text recognition, machine translation, and scene text synthesis and their task-specific design enhancements. (iii) Training a visual translation model with real-world images is challenging due to the lack of large-scale paired scene text images in different languages. Therefore, we use synthetic images for training. We present a method to generate paired images with words sharing the same visual properties, creating VT-Syn, a synthetic dataset of 600K paired visually diverse English-Hindi scene-text images. To evaluate performance on real images, we provide extensive annotations of translated text from three users. These benchmark datasets will support future research in visual translation. (iv) Due to the lack of principled evaluation metrics for visual translation tasks in the literature, we propose a set of automatic and user evaluation metrics. We believe these metrics will help track the progress of visual translation tasks effectively.
@@ 2 Related Work
Translation of Text in Images: Recent years have seen growing interest in translating text within images, both in research and commercial domains. Current works primarily focus on recognition and translation methods for scene text [20, 16], prioritizing accurate translation without addressing visually consistent text generation. A popular commercial product – Google lens  [^1] https://lens.google/#translate also falls into this category. These approaches often resort to simply overlaying translated text on source images. While some studies explore end-to-end methods for text translation in images that generate text directly in pixel space [22, 15], they typically deal with limited visual diversity in document-style images with plain backgrounds and fixed fonts – without tackling the complexities of scene text that we aim to address. The closest solution to our problem is Google Translate for images [^2] https://translate.google.com/?op=images, a commercial product for visual translation of diverse scene text. However, its underlying technology remains proprietary and closed-source. We emphasize the need for the research community to study this problem openly, establish proper open-source solutions, create a benchmark, and define evaluation criteria – goals we pursue in this paper. Moreover, we observe that Google Translate still lacks translation quality and often fails to produce visually consistent results for complex cases, underlining the potential for better approaches. In Section 6.2, we provide a qualitative comparison between our work and Google Translate.
Figure 2: Outline of proposed cascaded baseline for Visual Translation. We use state-of-the-art approaches for scene text recognition, machine translation, and scene text synthesis to design variants of our baseline. Moreover, we further investigate the scene-text synthesis and propose an extension to existing SRNet architecture.
Editing Text in Images: The problem of editing text in images has witnessed significant research interest in recent years [31, 32, 28, 18, 14, 27]. This task aims to modify scene text to target content while retaining visual properties of the original text. SRNet [31] is one such method that learns an end-to-end GAN-based style-retention network. SwapText [32] improved upon the SRNet architecture by modeling the geometrical transformation of the text using spatial points. More recently, TextStyleBrush [14], RewriteNet [18], and MOSTEL [27] introduce a self-supervised training approach on real-world images for this task. Further, TextStyleBrush is evaluated on handwritten images as well. Authors in [28] proposed a character-wise text editor model for this task. However, their approach assumes source and target text instances are of the same length, which is not always true, especially in the translation task. A more recent approach, MOSTEL [27], also introduces stroke-level modifications to produce more photo-realistic generations. Despite these advances, these methods only address the cross-lingual editing problem, which is just a component of the visual translation process and is insufficient on its own for achieving visual translation. Our work aims to address the task of visual translation and its complexities more comprehensively.
@@ 3 Proposed Visual Translation Baseline
The task of Visual Translation can be reduced to a sequence of sub-tasks: locating and reading text in scene images, translating the text into the target language, and generating the final image containing the translated scene text. Motivated by this observation, we propose a cascaded approach to visual translation by combining models for (i) scene text recognition, (ii) machine translation, (iii) scene-text synthesis, and (iv) seamlessly blending the generated scene text into the image. These sub-tasks are well-explored independently in computer vision literature; thus, we benefit from the availability of trained models. Further, such an approach can perform generation at the word or phrase level, which can help preserve the consistency of non-text regions in the image. The outline of our cascaded baseline is illustrated in Fig. 2 and a detailed illustration is provided in Fig. 3. We describe each module in detail in the following sections.
Training and evaluation of a visual translation baseline require real-world images in the form of $(I,I^{\prime})$ where $I$ and $I^{\prime}$ are visually identical images containing corresponding scene text in two different languages with matching font and style. However, such instances are not easily available in the real world. We mitigate this data scarcity challenge by directly synthesizing the desired data: we generate paired scene-text images that are (i) identical in the image background and (ii) matching in font and style. A few examples are shown in Fig. 4.
In generating synthetic samples, we use a large corpus of words in both languages, as well as a diverse collection of fonts. To simulate real-world scene text, we also render the images on natural backgrounds, as well as vary the orientation, positioning, and size of the scene text in images. A more detailed procedure for generating the synthetic data is provided in Section 4.1.
@@
(ii) Machine Translation of Text. After obtaining the recognized text in the source language $L$, we map each instance to the desired target language $L^{\prime}$ using an off-the-shelf neural machine translation method. We test our model with two state-of-the-art neural machine translation methods, namely IndicTrans2 [8] and M2M100 [7]. IndicTrans2 is trained on a large collection of Indic languages (including Hindi), whereas M2M100 is a more general translation method trained on a diverse collection of languages with support for Indic languages as well.
Figure 3: Our proposed baseline extends the SRNet scene text synthesis approach by decoupling background and foreground generation. More details provided in Section 2.
@@ 3.1 Design Enhancements
To enhance the design, a series of refined and newly introduced steps have been implemented. The process begins with the detection and recognition of text, after which numbers, websites, and email addresses are filtered out using regular expressions. Note that these elements do not need to be translated. Words are then grouped into paragraphs and lines based on the geometry and coordinates of the bounding box in conjunction with a heuristic function. These paragraphs are translated and segmented into lines, ensuring alignment with the proportion of lines present in each original paragraph. Through cubic spline interpolation, new coordinates for each word within a line are determined, which are then linked back to the original crops of the words. Depending on the new width of the translated words, adjustments are made to the crops—either cutting or replicating them—to maintain the original style of the text.
The process is finalized by accurately positioning the new words on the image using the developed method. Although heuristically designed, this step shows a significant boost in translation quality, as shown in the experiments.
@@ 3.2 Baseline Variants
Figure 4: VT-Syn dataset examples, which contains paired Eng $\rightarrow$ Hin and Hin $\rightarrow$ Eng images with diverse fonts, text colors, sizes, orientations, and background images of natural scenes, textures, and plain colors.
@@ 4.1 VT-Syn: Synthetic Training Data
For training the scene text synthesis components of our pipeline, we need paired images of text in different languages with identical visual properties (style, font, orientation, and size.). It is extremely difficult to get visually identical scene images with text in different languages in the real world, and it is even more difficult to generate accurate skeleton images required for training SRNet, MOSTEL, as well as our proposed SRNet++ method. Thus, we rely on generating highly diverse synthetic images. We introduce VT-Syn, a synthetically generated corpus of 600K visually diverse paired bilingual word images in pairs of English-Hindi as well as Hindi-English.
We utilized an Indic-language scene-text image generator [23] and modified it to generate samples of scene-text in paired languages with controllable parameters for font, style, color, and spatial transformations to ensure visual diversity. Each sample contains a source image, a target word image, a background image, a foreground image, and a target image. We also generate source word images, source and target masks, and skeleton images based on the requirements of various scene-text synthesis architectures. We collect 291 publicly available fonts that support both Roman and Devanagari scripts and use a vocabulary of 3K commonly used words in both languages.
A few samples from VT-Syn are shown in Fig. 4. Note that the paired image words do not have to be translations as the STS module has to particularly learn to render the target word using the same style as the source image.
@@ 4.2 VT-Real: Real Test Dataset
For the purpose of evaluation, we propose VT-Real, which contains images from ICDAR 2013 [13] and Bharat Scene Text Dataset [1] to evaluate English-to-Hindi and Hindi-to-English translations, respectively. We filter images of moderate complexity [^3] As a first work on Hindi-to-English and English-to-Hindi translation, we have opted not to include highly complex curved and occluded text. from these two sources. In all, our dataset contains 269 images and 1021 words. These images were given to three human annotators to translate the text from Hindi to English and vice versa. A few example translation annotations of this dataset are shown in Fig. 5. Even though the above-mentioned datasets have no ground truths for Visual Translation (i.e., scene text in the target language), they are still useful for automatic evaluation proposed in the next section.
Figure 5: A few examples from VT-Real dataset, showing image and Eng-Hin and Hin-Eng ground truth translations, manually annotated by three independent annotators (referred to as users here).
@@ 5 Performance Metrics
Evaluating visual translation methods is complex, even more so than evaluating machine translation. While the evaluation of machine translation has been a longstanding research area in the NLP community, recent research has saturated the use of metrics such as BLEU, METOER, and ROUGE; visual translation poses several additional challenges. Unlike machine translation, which is typically evaluated for a sentence or paragraph of text, in visual translation, one has to evaluate the correctness of translation for a single word or a small set of words or phrases. Further, It requires not only assessing the linguistic accuracy of the translation but also ensuring the preservation of background and font properties.
@@ 5.1 Automatic Evaluation
(iii) $\mathbf{VT\text{-}score}$: For high-quality visual translation, it is important to have a high BLEU score, and perception quality along with font style preservation is required. Due to the absence of a robust automatic model that can verify cross-lingual font style similarity, we only consider Translation Quality and perception quality to compute combined $vt\text{-}score$ as follows:
$vt\text{-}score=\frac{2\cdot TQ\cdot PQ}{TQ+PQ}.$ (1)
Please note that for images that contain only one word, we employ BLEU-1 instead of BLEU-2 to assess translation quality. For the remaining images, BLEU-2 is utilized in the above mentioned scoring measure.
@@ 5.2 User Evaluation
Despite the availability of automatic evaluation metrics, as discussed above, user evaluation is crucial for assessing the accuracy, usability, and effectiveness of visual translations. User feedback is essential for evaluating the clarity, cultural appropriateness, and accessibility of translations. To this end, we conducted an extensive user evaluation with four human users aged 20 to 25 who hold graduate degrees and are proficient in both Hindi and English. They reviewed each visual translation baseline using Beamer slides: slides for metrics (ii) and (iii) featured single output images, while those for metrics (i) and (iv) displayed both input and output images together. The user evaluation metrics are described here:
@@
(i) Translation Quality (TQ) (score range: 1-4): This criterion focuses on the accuracy of the translation. Users were asked to rate whether the translated text accurately conveys the meaning of the original text. A higher score indicates a more accurate translation. The different ratings by users convey the following: 4: Linguistically and culturally totally correct translation. 3: Some words are correct; translation can be improved. 2: Very few words are correct, and significant improvement is required. 1: Totally incorrect translation.
@@
(ii) Readability (R) (score range: 1-4): This criterion evaluates how easily the translated text can be read within the scene image. Factors such as font size, contrast, and placement of the text may influence readability. A higher score indicates better readability. The different readability ratings by users convey the following: 4: Clearly readable. 3: Can read with some effort. 2: Can read with significant effort; some words are not readable. 1: No text present in the target language.
(iii) Perceptual Quality (PQ) (score range: 1-4): This criterion assesses how well the translated text blends into the scene image, making it difficult to distinguish from a real image. A higher score indicates better integration of the translated text with the scene. Users were asked to rate approaches based on the following: 4: Very clear, looks like real image. 3: Clear image, but some patches are present if carefully seen. 2: There are a lot of patchy effects; looks like a fake image. 1: Too much patchy effect; for sure, it is a fake image.
@@
(iv) Source Style Preservation (SSP) (score range: 1-4): This criterion examines whether the translated text preserves the style, font, color, and other visual attributes of the original text in the scene image. A higher score indicates that the translated text maintains consistency with the source text in terms of visual presentation. 4: Font style, size, color, and background are coherent to the source. 3: Only 2 or 3 of the following: font style, size, color, and background are coherent to the source. 2: Only 1 or 2 of the following: font style, size, color, and background are coherent to the source. 1: No source-style preservation.
Method STR MT STS D.E. TQ (BL-1) TQ (BL-2) PQ VT-score
@@
B-4 DBNet+ParSeq M2M SRNet++ 03.22 02.19 55.60 03.81
B-3 Oracle M2M SRNet++ 04.20 02.89 55.58 04.97
B-2 Oracle M2M Mostel 02.03 01.40 53.41 02.46
B-1 Oracle M2M SRNet 04.20 02.86 53.82 04.92
Table 1: Automatic Evaluation to evaluate baselines for visual translation. We report translation quality (TQ) using BLEU-1 (BL-1) and BLEU-2 (BL-2) metrics and perception quality (PQ). D.E.: Design Enhancements. More details in Section 5.1.
6 Experiments
In this section, we comprehensively evaluate scene-text to scene-text translation baseline approaches discussed in Section 3.2 using both automatic and user evaluation metrics proposed in Section 5. We use VT-Real introduced in Section 4 for all our evaluation.
The automatic evaluation results are reported in Table 1. We observe that SRNet++ clearly emerges as the best scene text synthesis approach as compared to other existing architectures. The proposed design enhancements also significantly boost translation quality while maintaining nearly identical perceptual quality. We also observe that usage of IndicTrans2 as a translator consistently leads to an increase in translation quality. The state-of-the-art scene text recognition approaches are as good as ground truth annotations (Oracle) in the case of detecting and recognizing English text.
We further perform a rigorous user study using metrics presented in Section 5.2. As discussed in this section, we have collected user feedback from four qualified users and report mean scores of TQ, R, PQ, and SSP in Table 2. These scores nearly align with observations made via automatic evaluation. We also observe that there is significant room for improvement on all these metrics, indicating the challenge associated with the task.
Method  STR MT STS D.E. TQ R PQ SSP